The Sisyphean Company has a bond outstanding with a face value of $1,000 that reaches maturity in 15 years. The bond certificate indicates that the stated coupon rate for this bond is 8.8% and that the coupon payments are to be made semiannually.

Assuming the appropriate YTM on the Sisyphean bond is 10.8% , then the price that this bond trades for will be closest to:

a. $853

b. $1,194

c. $682

d. $1,024
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 15 * 2 = 30 semi-annual periods, C = 8.8% * $1000/2 = $44 semi-annually, i = 10.8%/2 = 5.4% (semi-annually)

1- P = 44* - (1+0.054)30 0.054 1000 (1 + 0.054)30

P = $646.61 + $206.43

P = $853.04

P = $853 (Option a)